Question: Devon purchased tickets to an air show for 9 adults and 2 children. The
total cost was $252. The cost of a childʹs ticket was $6 less than the cost of
an adultʹs ticket. Find the price of an adultʹs ticket and a childʹs ticket.
What i came up with:
I tried to to get what x and why are.
i put : x= 9 adults y= 2 children.
i know that child ticket is $6 less then which me the equation has subtraction.
i thought it might be 9x-2y=252 but it doesnt make sense.
please send step by step so i can understand how to do this correctly. thank you


Answer by lwsshak3(11628)

Question: Devon purchased tickets to an air show for 9 adults and 2 children. The
total cost was $252. The cost of a childʹs ticket was $6 less than the cost of
an adultʹs ticket. Find the price of an adultʹs ticket and a childʹs ticket.
***
let x=cost of adult ticket
x-6=cost of child's ticket
9x+2(x-6)=252
9x+2x-12=252
11x=264
x=24
x-6=18
cost of adult ticket=$24
cost of child's ticket=$18